### v1

# Hybrid-Emba3D: Geometry-Aware and Cross-Path Feature Hybrid Enhanced State Space Model for Point Cloud Classification

Abstract
The point cloud classification tasks face the dual challenge of efficiently extracting local geometric features while maintaining model complexity. The Mamba architecture utilizes the linear complexity advantage of state space models (SSMs) to overcome the computational bottleneck of Transformers while balancing global modeling capabilities. However, the inherent contradiction between its unidirectional dependency and the unordered nature of point clouds impedes modeling spatial correlation in local neighborhoods, thus constraining geometric feature extraction. This paper proposes Hybrid-Emba3D, a bidirectional Mamba model enhanced by geometry-feature coupling and cross-path feature hybridization. The Local geometric pooling with geometry-feature coupling mechanism significantly enhances local feature discriminative power via coordinated propagation and dynamic aggregation of geometric information between local center points and their neighborhoods, without introducing additional parameters. The designed Collaborative feature enhancer adopts dual-path hybridization, effectively handling local mutations and sparse key signals, breaking through the limitations of traditional SSM long-range modeling. Experimental results demonstrate that the proposed model achieves a new SOTA classification accuracy of 95.99% on ModelNet40 with only 0.03M additional. Our code and weights are available at https://github.com/L1277471578/Hybrid-Emba3D.
keywords:
State Space Model, Point Cloud, Classification, Local geometry
journal: Arxiv
[…]
[label1] organization=Xi’an Technological University,city=Xi’an, postcode=710021, state=Shaanxi, country=China
[…]
[label2]organization=Changchun University of Science and Technology, city=Changchun, postcode=130022, state=Jilin, country=China
1 Introduction
Point clouds directly represent the geometric features of objects through 3D coordinates, compensating for the lack of spatial information in 2D data and providing basic data support for perception tasks such as 3D object recognition and scene understanding[19]. However, the inherently disordered nature of point cloud data necessitates permutation invariant processing, and effective extraction of local geometric features is constrained by nonuniform distribution, these intrinsic characteristics present challenges for improving the accuracy and generalization capability of 3D point cloud classification models.
In response to the challenges of point cloud disorder and local feature extraction, researchers have explored. Among them, PointNet[4] innovatively integrates multi-layer perceptrons and symmetric functions, and builds a processing framework with permutation invariance for the first time, laying the foundation for the direct representation of the disordered features of point clouds. Subsequent research further captures local geometric features by developing point cloud convolution operations[44, 20, 47]. However, the convolution-based method is essentially local feature aggregation, which makes it difficult to model the long-distance dependency in point cloud data and limits its performance improvement in complex scenes. Some researchers have shifted their focus to Transformer[38] architectures with global modeling capabilities to overcome the limitations of the convolution method.
The PointTransformer[8] pioneered the application of self-attention mechanisms in 3D point cloud processing, utilizing positional encoding and multi-head attention mechanisms to establish comprehensive modeling of global dependencies within point clouds. While subsequent studies have enhanced feature representation capabilities through refined local regional modeling and hierarchical architectural designs, the intrinsic quadratic computational complexity of Transformer-based frameworks persistsently suffers from dual constraints on computational resources and inference speed[18].
State Space Model (SSM) is an emerging sequence modeling method in the field of deep learning. It realizes efficient long sequence modeling with the mathematical characteristics of continuous state representation and linear time-invariant systems. By introducing a selective mechanism with input dependency, the Mamba model[11] innovatively constructs a state space architecture with conditional computing characteristics. It combines the timing dependence modeling capabilities of RNNs[22] with the parallel computing advantages of the Transformer to achieve linear time complexity of long-distance dependence modeling. However, the irregular topological characteristics of point cloud data make it difficult for the SSM discrete state transfer mechanism to effectively capture geometric correlations. Furthermore, SSM’s unidirectional scanning paradigm is inadequate in modeling the intrinsically bidirectional spatial dependencies embedded within point clouds, particularly in local feature refinement processes where this limitation inevitably induces structural information loss of geometric details.
To address these limitations, this paper proposes a Geometry-Aware and Cross-Path Feature Hybrid Enhanced Mamba Model (Hybrid-Emba3D), which realizes lightweight local spatial information feature enhancement and feature cross-path fusion, improving the ability to extract local and global information.
The main contributions of this work can be summarized as follows:
  • 1.
    A lightweight feature coupling mechanism based on geometric priors is proposed, which couples geometric feature information between center point features and neighboring point spatial geometric information, enriching local geometric representations.
  • 2.
    Designing a multi-dimensional hybrid feature enhancement module with channel-wise and multi-scale learning capabilities, which enhances context-aware modeling capacity for SSM.
  • 3.
    The proposed model demonstrates exceptional performance, particularly outperforming existing SOTA models on the ModelNet40 benchmark and establishing new performance benchmarks. As shown in Figure 1.
Figure 1: Comprehensive comparison of Hybrid-Emba3D with other SOTA models
2 Related work
2.1 Deep learning for point cloud data
Effective point cloud classification requires the collaborative modeling of local geometric details and global spatial context. Maturana[29] et al. proposed VoxNet, which organizes unstructured point clouds into 3D voxel grids and applies conventional 3D convolutional neural networks to capture comprehensive spatial contextual features. Based on PointNet, PointNet++[31] employs Farthest Point Sampling (FPS) to construct hierarchical architectures while enhancing localized detail extraction through its local feature aggregation mechanism. Its density-adaptive modules effectively mitigate multi-scale learning interference caused by non-uniform point density distributions. Subsequent approaches encompass point-based direct convolution[36, 26, 9, 42], Volume-Based convolution [34, 35, 17, 43], and graph-based convolution[41, 28, 40, 24]. They are all attempts to enhance local feature extraction through spatial transformations or deformable convolutional kernels. However, these methodologies fundamentally suffer from the mismatch between fixed receptive fields and irregular point distributions.
In 2021, PointTransformer[8] was the first to introduce the Transformer architecture, marking a new stage in global context modeling for point cloud processing. Based on relative position coding, its self-attention mechanism can adaptively establish geometric correlations between point cloud elements and significantly improve the scene semantic understanding ability through global dependency modeling[30, 48, 51, 14, 45]. Notably, the architecture’s intrinsic permutation equivariance naturally conforms with the unordered nature of point clouds, eliminating the need for data rearrangement or interpolation required by voxel/grid-based methods, thereby avoiding the information loss caused by three-dimensional coordinate discretization.
The conflict exists between the sparse distribution characteristics of point cloud data and the dense connectivity assumption underlying standard attention mechanisms, leading to redundant attention weight allocations. Researchers have developed sparse attention strategies to improve this. Such as DSVT[10] with hierarchical feature aggregation demonstrates optimized computation-accuracy balance through adaptive sparse attention patterns; Point-BERT[48] self-supervised pretraining via masked point modeling, effectively reducing annotation dependency; PatchFormer[50] significantly reduces computational complexity through geometrically consistent patch partitioning and streamlined patch-level attention mechanisms.
While convolution-based architectures exhibit linear computational complexity, their localized receptive fields fundamentally constrain long-range contextual modeling capabilities. Conversely, Transformer models achieve global interaction through self-attention mechanisms, yet are inherently limited by quadratic computational complexity. In this context, state space models (SSMs) like Mamba demonstrate potential for point cloud processing, achieving linear modeling characteristics in long-sequence modeling that significantly reduce both computational and spatial overhead while preserving global dependency learning.
2.2 State Space Model
SSM, a classic method in control theory, achieves dynamic system modeling by constructing the state space. As control theory evolved, the limitations of traditional input-output models have gradually emerged. SSM establishes a system state space based on input signal characteristic variables, uses state variables to characterize the dynamic evolution of the system fully, and abstracts the system into a mathematical model containing state and output equations. The SSM structure are shown in Figure 2.
Figure 2: Mamba block and SSM structure
Discretize the continuous-time linear time-invariant (LTI) state space model using the zero-order hold (ZOH) method. The discretized equation of state is given in equation 1.
${}\left\{\begin{matrix}{{h}_{k}}&=\bar{A}\cdot{{h}_{k-1}}+\bar{B}\cdot{{x}_{k}}\\ {{y}_{k}}&=\bar{C}\cdot{{h}_{k}}+\bar{D}\cdot{{x}_{k}}\\ \end{matrix}\right.$ (1)
Where: $\bar{A}$ is a discrete state transition matrix, $\bar{B}$ is a discrete control input matrix, $\bar{C}$ is a discrete observation matrix, and $\bar{D}$ is a discrete feedforward matrix. ${h}_{k}$ is the current state variable, ${h}_{k-1}$ represents the hidden state of the system at the discrete time point $k$-1, containing historical information up to that time, and $k$ is the discrete time step index.
Traditional SSMs[12, 13, 2]employ static predefined parameters for matrices $\bar{A}$ , $\bar{B}$ , and $\bar{C}$ , resulting in rigid information propagation pathways. In contrast, Mamba’s SSM[11] layer implements a selective mechanism that dynamically generates timestep-specific parameters $\Delta$ conditioned on input data. This enables adaptive modulation of the discretization processes for B and C matrices, thereby establishing context-sensitive information propagation paths. Generate parameters $\Delta$, $\bar{B}$, $\bar{C}$ in real time from the input sequence through linear transformation:
${{B}_{k}}=\text{Linea}{{\text{r}}_{B}}({{x}_{k}})\in{{\mathbb{R}}^{N\times 1}}$ (2)
${{C}_{k}}=\text{Linea}{{\text{r}}_{C}}({{x}_{k}})\in{{\mathbb{R}}^{N\times 1}}$ (3)
${{\Delta}_{k}}=\text{softplus}(\text{Linea}{{\text{r}}_{\Delta}}({{x}_{k}}))\in{{\mathbb{R}}^{+}}$ (4)
Where: Linear represents linear projection, and $\Delta$ is the step length at which ZOH discretizes the continuous equation.
Dynamically adjust the discretization process according to the input:
${{{\bar{A}}}_{k}}=\exp({{\Delta}_{k}}{{A}_{k}})$ (5)
${{{\bar{B}}}_{k}}={{({{\Delta}_{k}}{{A}_{k}})}^{-1}}(\exp({{\Delta}_{k}}{{A}_{k}})-I){{\Delta}_{k}}{{B}_{k}}$ (6)
${{{\bar{C}}}_{k}}={{C}_{k}}$ (7)
Plugging parameter $\bar{A}$, $\bar{B}$, $\bar{C}$ into Equation 1, the selective SSM still follows the LTI property in the continuous time domain, thus allowing efficient GPU-accelerated convolution computing of the discretized state space equation:
$\bar{K}=(\overline{{{C}_{k}}{{B}_{k}}},{{\overline{{{C}_{k}}{{A}_{k}}B}}_{k}},...,\overline{{{C}_{k}}A}_{k}^{L-1}{{\overline{B}}_{k}})$ (8)
${{y}_{k}}={{x}_{k}}*\bar{K}$ (9)
Where: $\bar{K}$ is the global convolution kernel, $L$ is the length of the convolution kernel.
In summary, the selective SSM’s continuous-time equation is formally consistent with traditional SSM, but content-aware modeling is realized through the dynamic parameterization of the discretization process. This design paradigm transcends LTI limitations while inheriting the theory of continuous systems, enabling Mamba to simultaneously attain computational efficiency comparable to RNNs (O(n) complexity) and expressive power equivalent to Transformer architectures.
PointMamba[21] pioneers the introduction of Mamba architecture into point cloud processing, which significantly improves efficiency while maintaining high accuracy, demonstrating the potential of Mamba in 3D vision tasks.
3 Hybrid-Emba3D
While Mamba achieves input-dependent parameter adaptation through selective gating mechanisms, its core architecture remains anchored in RNN-style sequential recurrence. This design introduces inherent constraints in global context sensitivity during long-range dependency modeling, potentially limiting the comprehensive capture of complex interaction patterns.
3.1 Overview
This work establishes a hybrid framework that effectively enhances the representation performance of the Mamba. It integrates the Local geometric pooling (LGP) with geometric feature coupling and a bidirectional SSM enhanced by the collaborative feature enhancer (CoFE-BiSSM). The architecture of the model is shown in Figure 3.
Figure 3: The architecture of Hybrid-Emba3D. The input point cloud is sampled into local patches via FPS and KNN, and the coordinates and features are aggregated into embedding vectors through lightweight PointNet. After incorporating positional encoding and classification tokens, input the Hybrid-Emba3D encoder. The encoder core is composed of alternately stacked CoFE-BiSSM and LGP, supplemented by normalization and residual connections. Task-specific heads finally decode the transformed features for the classification of objects.
Given an input point cloud $P\in{{\mathbb{R}}^{N\times 3}}$ with $N$ points, the framework first initializes $L$ centroid points via Farthest Point Sampling (FPS) to form a set ${{P}_{\text{center}}}\in{{\mathbb{R}}^{L\times 3}}$. For each centroid ${{P}_{\text{center}}}$, the method retrieves its K-Nearest Neighbors (KNN) from the original cloud to construct local patches $x_{p}^{i}\in{{\mathbb{R}}^{K\times 3}}$ serving as structural spatial primitives. A lightweight PointNet variant subsequently performs hierarchical feature encoding on each patch, with the resultant local descriptors simultaneously serving as input token sequences for Hybrid-Emba3D architectures.
Adhering to the ViT’s[7] work, the architecture embeds a learnable [class] token preceding $L$ local tokens for global representation aggregation. Each patch ${{x}_{p}}\in{{\mathbb{R}}^{L\times K\times 3}}$ is projected and mapped into a C-dimensional feature vector ${{W}_{patch}}({{x}_{p}})\in{{R}^{L\times C}}$ as its initial patch embeddings. The [class] token is vertically stacked with all point cloud block features, forming a sequence matrix with a shape of ($L$+1)×$C$ ($L$ patch features $+$ one [class]). After superimposing position encoding ${P}_{pos}$, the geometric structure information of the original point cloud is preserved through patch position encoding, providing an initial representation for subsequent hierarchical encoding.
${{Z}^{0}}=[{{x}_{\text{cls}}}\text{+}P_{\text{pos}}^{(0)},{{W}_{patch}}(x_{p}^{1})+P_{\text{pos}}^{(1)},\cdots,{{W}_{patch}}(x_{p}^{L})+P_{\text{pos}}^{(L)}]$ (10)
Where: ${Z}^{0}$ is the initial input layer feature, ${{x}_{cls}}\in{{\mathbb{R}}^{1\times C}}$ is the [class] token, ${{W}_{patch}}(x_{p}^{1}),{{W}_{patch}}(x_{p}^{2})\cdots,{{W}_{patch}}(x_{p}^{L})$ is the concatenated point cloud patch feature processed by the lightweight PointNet variant (${W}_{patch}$), and ${P}_{pos}$ is the positional encoding.
The encoder employs dual-phase processing (The encoder block is shown in Figure 3 (b)): For the $i$-th layer, input features ${Z}^{i-1}$ undergo normalization before the LGP component executes geometric-aware local feature extraction to generate ${{\hat{Z}}^{i}}$. Post residual summation, secondary normalization precedes the CoFE-BiSSM component, which supplants conventional attention mechanisms, enabling dynamic feature enhancement and long-range dependency capture, ultimately outputting the current layer representation ${Z}^{i}$. The overall structure retains the residual connection. The process is as follows:
${{{\hat{Z}}}^{i}}=\text{LGP}(\text{Norm}({{Z}^{i-1}}+{{P}_{\text{pos}}}))+{{Z}^{i-1}},$ (11)
${{{Z}}}^{i}=\text{CoFE-BiSSM}(\text{Norm}({{{\hat{Z}}}^{i}}))+{{{\hat{Z}}}^{i}}$ (12)
The architecture employs task-specific heads (e.g., classification MLPs) for diverse downstream tasks, while integrating learnable positional embeddings within each encoder layer to enhance spatial awareness. This structure ingeniously transplants the achievements of the Transformer into the Mamba series.
3.2 Local geometric pooling with geometry-feature coupling mechanism
Local features are crucial for point cloud feature learning. The local features in point clouds are usually obtained by constructing a local domain using KNN and then performing feature fusion. Mamba3D simplifies local feature extraction to feature propagation and aggregation operations, but ignores the spatial geometric information of local neighborhoods. Therefore, we have designed a local geometry-feature coupling mechanism, which mainly contains local feature normalization, local geometry patch normalization, and feature coupling operations.
Figure 4: Illustration of LGP. This process selects neighboring points through FPS and KNN, extracts normalized features from the center points, and element-wise multiplication with the broadcast extended local geometric weights to achieve the fusion of features and geometric information, followed by a shared MLP for advanced feature extraction.
According to Section 3.1 and Figure 4, $L$ local domains with $k$ neighbors are constructed using KNN around each center point. The center point of these local domains is ${{P}_{\text{center}}}\in{{\mathbb{R}}^{L\times 3}}$ , and the coordinates of the neighborhood point are ${{k}_{i}}$.
For the local feature normalization, given the center point feature and its K nearest neighbor feature ${{\tilde{F}}_{K}}\in{{\mathbb{R}}^{L\times K\times C}}$, the relative feature offset $\Delta F={{F}_{K}}-{{F}_{C}}$ is first calculated and channel normalized:
${{\tilde{F}}_{K}}=\frac{\Delta F}{\sqrt{Var(\Delta F)+\varepsilon}}$ (13)
The normalized neighbor features and the center point features are concatenated along the channel, and the propagation features can be generated through learning an affine transformation:
${{\hat{F}}_{K}}\in{{\mathbb{R}}^{L\times K\times 2C}}=Concat({{\tilde{F}}_{K}},{{F}_{C}})\odot\gamma+\beta$ (14)
Where: $\gamma$ and $\beta$ are trainable proportional coefficients and shift vectors, respectively.
That implements feature-aware adaptive updates by explicitly encoding the center-neighborhood feature relationship.
For the local geometric patch normalization, firstly compute the mean coordinate $\mu$ of $M$ points within the patch, then convert the coordinates of each point into offsets $\Delta p={{p}_{j}}-u$ relative to the patch centroid to eliminate global translational effects. Subsequently, calculate the coordinate standard deviation $\sigma={\sqrt{\sum\nolimits_{x=1}^{M}{||\Delta{{p}_{j}}||_{2}^{2}+\varepsilon}}}/{M}\;$, and normalize the coordinates by $p_{j}^{{}^{\prime}}={\Delta{{p}_{j}}}/{\sigma}\;$ to eliminate local scale differences. This process ensures that the geometric encoding achieves translation-scale invariance. The process can be formalized as:
$p_{j}^{{}^{\prime}}=\frac{{{p}_{j}}-\mu}{\sqrt{\frac{1}{M}\sum\nolimits_{k=1}^{M}{||{{p}_{k}}-\mu||_{\text{2}}^{\text{2}}+\varepsilon}}}$ (15)
Where: ${{p}_{j}},{{p}_{k}}\in{{\mathbb{R}}^{3}}$ are 3D coordinates of the $j$-th or $k$-th point in the local geometric patch; $\mu$ is the Centroid of the geometric patch, $\mu={\sum\nolimits_{k=1}^{N}{{{p}_{k}}}}/{M}\;$; $j$, $k$ are point indices, where $j$ is used for calculating the offset of individual points, and $k$ is used for summation traversal over all points; $\varepsilon$ is the numerical stability term.
This standardization operation enables local geometric features to have translation invariance and scale invariance, significantly improving the model’s generalization ability to geometric deformations. Furthermore, Gaussian spatial weights are constructed based on the Euclidean distances between each point in the local patch and the center point to model the geometric structure correlation:
$\Delta{{p}_{i}}={{p}_{i}}-{{p}_{c}},$ (16)
${{w}_{i}}=\exp(-||\Delta{{p}_{i}}|{{|}_{2}}),$ (17)
$W=Broadcast({{w}_{i}})\in{{\mathbb{R}}^{2C\times L\times K}}$ (18)
${{F}_{weighted}}={{{\hat{F}}}_{k}}\odot W$ (19)
Where: ${{p}_{i}},{{p}_{c}}\in{{\mathbb{R}}^{3}}$ represent the coordinates of patch points and the center point; ${{w}_{i}}\in{{\mathbb{R}}^{L\times K\times 1}}$ is a weight coefficient based on distance attenuation, ${{\hat{F}}_{k}}$ is a normalized feature tensor, with a shape of $2C$×$L$×$K$; $\odot$ represents element-wise multiplication; $Broadcast(\cdot)$ extends the weight vector to align with the feature dimension.
The final feature ${{F}_{agg}}\in{{\mathbb{R}}^{2C\times L}}$ is obtained using the Softmax weighted average to achieve adaptive aggregation of neighborhood features:
${{F}_{agg}}=\frac{\sum\nolimits_{k=1}^{K}{{{F}_{weighted}}\centerdot\exp({{F}_{weighted}})}}{\sum\nolimits_{k=1}^{K}{\exp({{F}_{weighted}})}}$ (20)
Finally, channel alignment is completed through the shared MLP.
This component constructs parameter-free geometric weights using Gaussian kernel functions, leveraging distance-dependent decay characteristics between points to reinforce local structural awareness. These weights exhibit well-defined physical significance, being solely geometry-driven without extra parameters. Through a coupling mechanism between spatial distribution and feature representation, it effectively enhances sensitivity to local structures.
3.3 Bidirectional SSM with the Collaborative feature enhancer
While Mamba achieves theoretical benefits in long-sequence modeling through SSMs, its RNN-based recurrent architecture inherently suffers from constrained effective receptive fields due to error accumulation in sequential processing, exhibiting a deficiency in long-range dependency modeling efficiency compared to Transformer architectures.
The proposed Collaborative feature enhancer (CoFE) component adopts the dual-path architecture for long-range sequence modeling, enabling efficient global context capture via grouped multi-scale fusion. The structure is shown in Figure 5.
Figure 5: Details of Collaborative feature enhancer
Given the input tensor $X\in{{\mathbb{R}}^{B\times C\times L}}$ divided into $g$ groups along the channel dimension:
$X^{\prime}=\operatorname{Reshape}(X,[B\cdot g,\frac{C}{g},L])$ (21)
Construct parallel dual-path features—the 3×3 convolutional path and the gated normalized path to extract local details and global context:
${{X}_{1}}=GN(X^{\prime}\odot\sigma(Con{{v}_{1\times 1}}({{P}_{Avg}}(X^{\prime}))))\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (22)
${{X}_{2}}=Con{{v}_{3\times 3}}(X^{\prime})\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (23)
Where: $GN$ means GroupNorm normalization, ${P}_{avg}$ is adaptive average pooling.
Perform cross-channel global feature compression on two paths, compress into a single channel, preserve information in the spatial dimension $L$, and serve as the basis for spatial attention weights:
$\phi({{X}_{i}})={{F}_{soft}}({{P}_{avg}}({{X}_{i}}))\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (24)
Where: ${F}_{soft}$ is the row-wise Softmax.
Subsequently, cross-path bidirectional interaction is performed to multiply the compressed feature matrices $\phi({X}_{1})$ and $\phi({X}_{2})$:
${{X}_{1}}\to{{X}_{2}}:\phi({{X}_{1}}){{X}_{2}}\in{{\mathbb{R}}^{B\cdot g\times 1\times L}}$ (25)
${{X}_{2}}\to{{X}_{1}}:\phi({{X}_{2}}){{X}_{1}}\in{{\mathbb{R}}^{B\cdot g\times 1\times L}}$ (26)
The output at this time is a correlation matrix between spatial positions, which is used to characterize the degree of feature matching at different positions ($L$ dimensions). The correlation matrix between spatial positions is normalized and activated by the gating mechanism to achieve nonlinear scaling of the normalized weights:
$W=\sigma(\phi({{X}_{1}}){{X}_{2}}+\phi({{X}_{2}}){{X}_{1}})$ (27)
Where: $\sigma(\cdot)$ is the Sigmoid.
Finally, merge the weights and adjust the output shape:
$\text{output}^{\prime}=X^{\prime}\odot W\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (28)
$\text{output}=\operatorname{Reshape}(\text{output}^{\prime},(B,C,L))$ (29)
CoFE employs dual pathways to capture local-global contexts, where a dynamic weight matrix explicitly fuses cross-scale features via bidirectional interaction (${X}_{1}\rightleftarrows{X}_{2}$). This interaction computes spatial affinities to eliminate unidirectional bias and adaptively models non-uniform dependencies in long sequences, delivering a lightweight attention enhancement for BiSSM (see Figure 6).
Figure 6: The structure of CoFE-BiSSM
The unidirectional nature of Mamba restricts global dependency capture in point cloud processing. Unlike Vision Mamba’s[53] token-level horizontal flipping, a reverse path through vertical flipping of feature channels is constructed. The resulting BiSSM module is formulated as:
$F^{\prime}=\text{CoFE}(Con{{v}_{1\times 1}}(F))$ (30)
[…]
Where:$Fli{{p}_{C}}(\cdot)$ is the Channel Flip operation.
This approach eliminates reliance on point cloud token order, enabling the model to focus on inherent distribution patterns of feature vectors rather than artificially constructed sequential orders. While Mamba’s core mechanism captures long-range dependencies through parameterized state transition equations, it demonstrates insufficient sensitivity to local transient features. The CoFE component complements the SSM framework by directly modeling inter-position relationships. The ${X}_{2}$ pathway reduces the limitations of SSMs in capturing local details through localized context aggregation. Subsequent ${X}_{1}\rightleftarrows{X}_{2}$ interaction facilitates complementary integration of feature correlation and provides more discriminative input to SSMs, enabling SSMs to dynamically perceive global contextual patterns while effectively modeling complex sequential dependencies.
4 Experiments
In this section, we first introduced hyperparameter settings and performance evaluation metrics. Secondly, conduct ablation experiments to evaluate the effectiveness of the components. Finally, the performance on target classification and small sample learning tasks was demonstrated.
4.1 Implementation Details
We conducted experiments under two distinct training protocols: training from scratch and fine-tuning. The cross-entropy loss function was adopted as the optimization objective for both settings. All experiments were performed on one NVIDIA TITAN RTX GPU without mixed-precision acceleration.
The model employs 12-layer encoders, feature dimension of 384. Point clouds are partitioned into 128 groups (Group Size=32) during preprocessing. Training uses AdamW optimizer with initial learning rate of 5e-4, weight decay of 0.05, cosine scheduler (CosLR), drop path rate of 0.2, warmup epochs of 10, and batch size of 32 for 300 epochs. This experiment employs overall accuracy (OA) as the classification performance indicator. Parameters quantify model complexity, and FLOPs measure computational complexity.
4.2 Object classification
Datasets: ModelNet40[46], a widely adopted benchmark dataset for 3D point cloud processing, comprises 12,311 structurally aligned 3D CAD models spanning 40 common object categories. ScanObjectNN[37], a real-world 3D object classification dataset, includes three variants: OBJ_BG (with Background), OBJ_ONLY (Vanilla), and PB_T50_RS (Perturbed with 50% Translation, Rotation, and Scaling). These variants respectively simulate practical challenges, with PB_T50_RS being the most challenging variant. The dataset contains $\sim$15K objects across 15 indoor object categories.
Training Settings: For fine-tuning, the pre-training employs the Point-MAE framework with masked self-supervised learning on the ShapeNetCore[3] dataset. Unless explicitly specified, consistent training parameters are maintained between scratch training and fine-tuning. The Hybrid-Emba3D classification task header utilizes a 3-layer MLP classifier (256→256→$C$ channels). In the ModelNet40 experiment, adopting the standard division (9843 train/2468 test), sampling 1024 points, and using scale&translation augmentation, while ScanObjectNN followed the original train/test split, sampling 2048 points, and using the rotation augmentation. The experimental results are shown in Table 1.
Table 1: Experimental results of ModelNet40 and ScanObjectiNN. Compared with methods of different architectures: 3D understanding architectures, Transformer-based architectures, •Diffusion-driven frameworks, Mamba-based architectures. * denotes the application of voting strategy. The results of other models are sourced from publicly published papers.
Methods  Reference Pre-training strategy ModelNet40 1k P ScanObjectNN P (M) F (G)
OBJ_BG OBJ_ONL PB_T50_RS
Supervised Learning Only
PointNet[4] CVPR 17 × 89.2 73.3 79.2 68.0 3.5 0.5
PointNet++[31] NeurIPS 17 × 90.7 82.3 84.3 77.9 1.5 1.7
DGCNN[41] TOG 19 × 92.9 82.8 86.2 78.1 1.8 2.4
DRNet[33] WACV-21 × 93.1 - - 80.3 - -
MVTN[15] ICCV 21 × 93.8 92.6 92.3 82.8 11.2 43.7
PointMLP[27] ICLR 22 × 94.5 - - 85.4±0.3 12.6 31.4
PointNeXt[32] NeurIPS 22 × 92.9 - 87.7±0.4 1.4 3.6
Transformer[38] NeurIPS 17 × 94.1 79.86 80.55 77.24 22.1 4.8
PointConT[25] JAS 23 × 93.5 - - 90.30 - -
Pointmamba[21] NeurIPS24 × - 88.30 87.78 82.48 12.3 3.6
Mamba3D[16] ACM MM24 × 93.4 92.94 92.08 91.81 16.9 3.9
Hybrid-Emba3D × 94.0 93.63 92.25 91.05 16.96 3.95
With Self-supervised Pre-training
Transformer[38] NeurIPS 17 OcCo 92.1 84.85 85.54 78.79 22.1 4.8
MaskPoint[23] CVPR 22 MaskPoint 93.8 89.30 88.10 84.30 22.1 4.8
Point-BERT*[48] CVPR 22 IDPT 93.4 88.12 88.30 83.69 22.1+1.7 4.8
Point-MAE*[30] ECCV 22 IDPT 94.4 91.22 90.02 84.94 22.1+1.7 4.8
PointGPT-S*[5] NeurIPS 23 94.0 91.6 90.0 86.9 29.2 5.7
Point-FMAE*[49] AAAI 24 Point-M2AE 95.15 95.18 93.29 90.22 27.4 3.6
PointDif[52] CVPR 24 PointDif - 93.29 91.91 87.61 - -
Pointmamba[21] NeurIPS 24 Point-MAE 93.6 94.32 92.60 89.31 12.3 3.6
Mamba3D*[16] ACM MM24 Point-MAE 95.1 95.18 92.60 93.34 16.9 3.9
Hybrid-Emba3D Point-MAE 95.58+1.58 93.80+0.27 93.12+0.87 93.09+2.04 16.96 3.95
Hybrid-Emba3D* Point-MAE 95.99+1.99 95.18+1.38 93.80+1.55 94.34+3.29 16.96 3.95
Comparison with existing SOTA methods: Hybrid-Emba3D achieves SOTA performance across multiple benchmarks. Without pretraining, the model attains 94.0% accuracy on ModelNet40, surpassing the comparable Mamba3D by 0.6%, while achieving the highest OBJ_BG accuracy of 93.63%. When employing Point-MAE pretraining, Hybrid-Emba3D breaks the 95.58% accuracy barrier on ModelNet40, further reaching 95.99% with the voting strategy, establishing a new record for this benchmark. Hybrid-Emba3D outperforms SOTA Transformer counterparts (PointGPT-S, Point-FMAE) by 1.99% and 0.84% in accuracy, while reducing Parameters by 41.9% and 39.1%, respectively. Compared to similar Mamba-based models (Pointmamba, Mamba3D), it achieves 0.89% and 1.0% accuracy gains on ModelNet40 and ScanObjectNN-PB_T50_RS (the most challenging benchmark) with only exchanged 0.06M Parameters and 0.05G FLOPs.
Feature visualization:High-dimensional features are reduced via t-SNE to a 2D projection distribution to visualize the clustering effect of features. As shown in Figure 7. On ModelNet40, the feature distribution is relatively scattered, and there are ambiguous inter-class boundaries, indicating that the Hybrid Emba3D needs to improve its ability to distinguish specific categories. While OBJ_BG and PB_T50_RS demonstrate prominent clustering tendencies with minor inter-class overlaps, OBJ_ONLY achieves superior class separation. The model maintains robust clustering performance under PB_T50_RS interference scenarios, confirming its resilience to environmental disturbances.
Figure 7: Details of Collaborative feature enhancer
Figure 8: Details of Collaborative feature enhancer
Analysis of analysis category confusion: Overall performance is good on ModelNet40 (see Table 1 and Figure 8), but similar geometric structures and overlapping local features result in a specific category-15 (dresser) accuracy of only 35% (consistent with t-SNE), and its confusion rates with category-26 (piano) and category-37 (tv_stand) are 35% and 25%, respectively.
Comparing OBJ_ONLY and OBJ_BG, the presence of background increases the misjudgment rate of locally structurally similar objects. For example, the accuracy of category-0 (bag) decreases from 100% to 88%, and the confusion rate of category-5 (chair) and 9 (shelf) reaches 13%. Conversely, it may also reduce the misclassifications rate caused by feature overlap, such as the accuracy rate of category-10 and 12 increased from 86% and 75% to 95% and 88%, respectively, and the misjudgment rate of category-9 (shelf) and 10 (table), category-11 (bed) and category-12 (pillow) drops from 9% and 12% to 0%. Under interference (PB_T50_RS), the misclassification distribution has shifted, meaning misjudged samples are redistributed among error categories while maintaining a stable overall misjudgment rate.
Table 2: Ablation studies on ModelNet40 and ScanObjectNN.
LGP CoFE-BiSSM ModelNet40 1k P ScanObjectNN P (M) F (G)
OBJ_BG OBJ_ONL PB_T50_RS
- - 93.39 91.60 89.84 90.32 16.93 3.86
- 93.48 92.57 90.88 90.67 16.93 3.86
- 93.56 92.05 90.02 90.49 16.96 3.95
94.00 93.63 92.25 91.05 16.96 3.95
Table 3: The experimental results of ModelNetFewShot. The results of other models are sourced from publicly published papers.
Methods  Reference 5-Way 10-Way
10-Shot 20-Shot 10-Shot 20-Shot
Supervised Learning Only
PointNet[4] CVPR 17 52.0±3.8 57.8±4.9 46.6±4.3 35.2±4.8
Transformer[38] NeurIPS 17 87.8±5.2 93.3±4.3 84.6±5.5 89.4±6.3
DGCNN[41] TOG 19 31.6±2.8 40.8±4.6 19.9±2.1 16.9±1.5
DGCNN+CrossPoint[1] CVPR 2022 92.5±3.0 94.9±2.1 83.6±5.3 87.9±4.2
Hybrid-Emba3D 90.5±3.8 96.0±3.2 86.3±5.1 92.0±3.4
With Self-supervised Pre-training
DGCNN+OcCo NeurIPS 17 90.6±2.8 92.5±1.9 82.9±1.3 86.5±2.2
OcCo[39] NeurIPS 17 94.0±3.6 95.9±2.7 89.4±5.1 92.4±4.6
ACT[6] NeurIPS 17 96.8±2.3 98.0±1.4 93.3±4.0 95.6±2.8
MaskPoint[23] CVPR 22 95.0±3.7 97.2±1.7 91.4±4.0 93.4±3.5
Point-BERT[48] CVPR 22 94.6±3.1 96.3±2.7 91.0±5.4 92.7±5.1
Point-MAE[30] ECCV 22 96.3±2.5 97.8±1.8 92.6±4.1 95.0±3.0
PointGPT-S[5] NeurIPS 23 96.8±2.0 98.6±1.1 92.6±4.6 95.2±3.4
Pointmamba[21] NeurIPS 24 96.9±2.3 99.0±1.1 93.0±4.0 95.6±2.8
Hybrid-Emba3D 93.3±5.0 98.9±1.3 90.4±5.7 95.6±3.9
4.3 Ablation study
To validate the complementary nature and distinct performance boundaries of core components LGP and CoFEBiSSM, across diverse scenarios, we conduct ablation studies on the Hybrid-Emba3D benchmarked on 3D point cloud classification tasks. The results as shown in Table 2
LGP: LGP is a core component focusing on local feature extraction. This component captures subtle geometric variations and topological structures on object surfaces by constructing non-linear transformation mechanisms within local neighborhoods. When solely activated, LGP achieves 94.00% accuracy on ModelNet40, demonstrating its strong geometric feature representation in ideal, background-free scenarios. Furthermore, it achieves accuracy gains of +0.97% (OBJ_BG), +1.04% (OBJ_ONLY), and +0.35% (PB_T50_RS) compared to the deactivated baseline, further validating its capability to resist background clutter and minor deformations through local geometric modeling. It is commendable that it introduces zero additional parameters (0M) and computational costs (0G FLOPs), highlighting superior lightweight design advantages.
CoFE-BiSSM: CoFE-BiSSM enhances long-range dependencies through multi-scale feature fusion and bidirectional modeling. It achieves 92.05% accuracy on OBJ_BG, slightly below LGP’s 92.57%, but with excellent efficiency (0.03M Parameters/0.09G FLOPs). In the ModelNet40 (93.56%) and OBJ-ONLY (90.02%) scenarios without background interference, the performance is relatively limited, while the rotation disturbance sensitivity on PB_T50_RS (90.32%) reveals insufficient local geometric modeling. Experiments demonstrate its effectiveness in global context reasoning for complex background tasks, but it requires integration with local feature extraction modules to address geometric sensitivity.
Joint effect: Component collaboration yields significant performance gains. At the peak of 94.00% (+0.61%) in ModelNet40, all subsets of ScanObjectNN surpassed the single component configuration comprehensively, with OBJ-BG improving by 2.03%, OBJ-ONLY improving by 2.41%, and PB_T50_S improving by 0.73%. They achieve feature complementarity with minimal overhead. In summary, LGP and CoFE-BiSSM focus on local geometric and global multi-scale modeling, respectively. While their strengths are scenario-specific, joint deployment enables complementary features and enhances generalization performance on different datasets.
4.4 Few-shot Learning
Datasets and Settings: ModelNetFewShot is the ModelNet40-based 3D few-shot benchmark for evaluating classification algorithms’ generalization with limited samples. It adopts N-way K-shot tasks, randomly selecting N classes (K training samples per class) and testing on the remaining instances. Results report mean accuracy ± standard deviation across 10 independent trials. The results are shown in Table 3.
Results: Without pretraining, Hybrid-Emba3D achieves 90.5% on 5-Way 10-Shot, surpassing Transformer (87.8±5.2) but lagging behind DGCNN-CrossPoint (92.5%). Scaling to 20-shot boosts accuracy by 6% (96.0%). For 10-Way tasks, it outperforms DGCNN-CrossPoint. With self-supervised pretraining, it is worth noting that Hybrid Emba3D outperforms PointGPT-S (98.6%) in 5-Way 20 Shot, achieving a quasi SOTA performance of 98.9%. However, the advantage has not been generalized to other settings. In the 5/10-Way task, the accuracy rate increased by 5.6% and 5.2% when the sample size increased from 10-Shot to 20-Shot, and the standard deviation dropped sharply by 74% and 32%, indicating that the model is sensitive to data sparsity, and data enhancement can significantly improve performance and stability.
5 Conclusion
In this paper, we present Hybrid-Emba3D, a novel mamba-based Architecture for point cloud classification task, designing two key components: the geometry-feature coupling mechanism that aggregates geometric information from neighboring points around each centroid and central feature to enhance local geometric representation, and A multidimensional feature hybrid enhancer employing dual-path Architecture to strengthen global context modeling for BiSSM. Notably, the total computational cost with the two components is very small. Especially in the operations of coupling local geometric information into central features in a parameter-free manner, incurring no additional parameters, that is completely driven by local neighborhoods’ intrinsic geometric relationships.
Extensive experiments demonstrate Hybrid-Emba3D’s superior performance across classification benchmarks, particularly achieving new SOTA accuracy on ModelNet40 at the cost of minimal parameters. These also show that the Mamba architecture has potential, yet insufficient exploration in point cloud processing.

### v2

@@
-Hybrid-Emba3D: Geometry-Aware and Cross-Path Feature Hybrid Enhanced State Space Model for Point Cloud Classification
+HyMamba: Mamba with Hybrid Geometry-Feature Coupling for Efficient Point Cloud Classification
 Abstract
-The point cloud classification tasks face the dual challenge of efficiently extracting local geometric features while maintaining model complexity.
-The Mamba architecture utilizes the linear complexity advantage of state space models (SSMs) to overcome the computational bottleneck of Transformers while balancing global modeling capabilities.
-However, the inherent contradiction between its unidirectional dependency and the unordered nature of point clouds impedes modeling spatial correlation in local neighborhoods, thus constraining geometric feature extraction.
-This paper proposes Hybrid-Emba3D, a bidirectional Mamba model enhanced by geometry-feature coupling and cross-path feature hybridization.
-The Local geometric pooling with geometry-feature coupling mechanism significantly enhances local feature discriminative power via coordinated propagation and dynamic aggregation of geometric information between local center points and their neighborhoods, without introducing additional parameters.
-The designed Collaborative feature enhancer adopts dual-path hybridization, effectively handling local mutations and sparse key signals, breaking through the limitations of traditional SSM long-range modeling.
-Experimental results demonstrate that the proposed model achieves a new SOTA classification accuracy of 95.99% on ModelNet40 with only 0.03M additional.
-Our code and weights are available at https://github.com/L1277471578/Hybrid-Emba3D.
+Point cloud classification is one of the essential technologies for achieving intelligent perception of 3D environments by machines, its core challenge is to efficiently extract local and global features.
+Mamba leverages state space models (SSMs) for global point cloud modeling.
+Although prior Mamba-based point cloud processing methods pay attention to the limitation of its flattened sequence modeling mechanism in fusing local and global features, the critical issue of weakened local geometric relevance caused by decoupling geometric structures and features in the input patches remains not fully revealed, and both jointly limit local feature extraction.
+Therefore, we propose HyMamba, a geometry and feature coupled Mamba framework featuring: (1) Geometry-Feature Coupled Pooling (GFCP), which achieves physically interpretable geometric information coupling by dynamically aggregating adjacent geometric information into local features; (2) Collaborative Feature Enhancer (CoFE), which enhances sparse signal capture through cross-path feature hybridization while effectively integrating global and local contexts.
+We conducted extensive experiments on ModelNet40 and ScanObjectNN datasets.
+The results demonstrate that the proposed model achieves superior classification performance, particularly on the ModelNet40, where it elevates accuracy to 95.99% with merely 0.03M additional parameters.
+Furthermore, it attains 98.9% accuracy on the ModelNetFewShot dataset, validating its robust generalization capabilities under sparse samples.
 keywords:
-State Space Model, Point Cloud, Classification, Local geometry
+Point cloud classification, State space model, Local geometry, Feature hybridization
 journal: ArXiv
 […]
-[label1] organization=Xi’an Technological University,city=Xi’an, postcode=710021, state=Shaanxi, country=China
+[label1] organization=School of Opto-electronical Engineering, Xi’an Technological University,city=Xi’an, postcode=710021, country=China
 […]
-[label2]organization=Changchun University of Science and Technology, city=Changchun, postcode=130022, state=Jilin, country=China
+[label2] organization=Xi’an Key Laboratory of Active Photoelectric Imaging Detection Technology, Xi’an Technological University, city=Xi’an, postcode=710021, country=China
 1 Introduction
 Point clouds directly represent the geometric features of objects through 3D coordinates, compensating for the lack of spatial information in 2D data and providing basic data support for perception tasks such as 3D object recognition and scene understanding [1].
-However, the inherently disordered nature of point cloud data necessitates permutation invariant processing, and effective extraction of local geometric features is constrained by nonuniform distribution, these intrinsic characteristics present challenges for improving the accuracy and generalization capability of 3D point cloud classification models.
-In response to the challenges of point cloud disorder and local feature extraction, researchers have explored.
-Among them, PointNet[4] innovatively integrates multi-layer perceptrons and symmetric functions, and builds a processing framework with permutation invariance for the first time, laying the foundation for the direct representation of the disordered features of point clouds.
-Subsequent research further captures local geometric features by developing point cloud convolution operations[44, 20, 47].
-However, the convolution-based method is essentially local feature aggregation, which makes it difficult to model the long-distance dependency in point cloud data and limits its performance improvement in complex scenes.
-Some researchers have shifted their focus to Transformer[38] architectures with global modeling capabilities to overcome the limitations of the convolution method.
-The PointTransformer[8] pioneered the application of self-attention mechanisms in 3D point cloud processing, utilizing positional encoding and multi-head attention mechanisms to establish comprehensive modeling of global dependencies within point clouds.
-While subsequent studies have enhanced feature representation capabilities through refined local regional modeling and hierarchical architectural designs, the intrinsic quadratic computational complexity of Transformer-based frameworks persistsently suffers from dual constraints on computational resources and inference speed[18].
-State Space Model (SSM) is an emerging sequence modeling method in the field of deep learning.
+However, the inherently disordered nature of point cloud data necessitates permutation invariant processing, and effective extraction of local geometric features is constrained by nonuniform distribution; these intrinsic characteristics present challenges for improving the accuracy and generalization capability of 3D point cloud classification models.
+Early approaches discretize unordered 3D point clouds into 3D voxel grids and employ 3D convolutional networks to capture spatial features.
+Subsequently, PointNet [2] innovatively builds a processing framework with permutation invariance, laying the foundation for the direct representation of the disordered features of point clouds.
+Follow-up research further captures local geometric features by developing point cloud convolution operations [3, 4, 5].
+Recent studies have introduced Transformer [6] architectures with global modeling capabilities, which utilize positional encoding and multi-head attention mechanisms to comprehensively model global dependencies within point clouds [7].
+Although subsequent research has enhanced feature representation through improved local region modeling and hierarchical architecture designs, they remain constrained by the inherent quadratic computational complexity of the Transformer framework [8].
+To circumvent the huge computational burden generated when establishing long-range dependencies, the state space model (SSM), as a novel approach for long-sequence modeling, has been introduced into point cloud learning.
 It realizes efficient long sequence modeling with the mathematical characteristics of continuous state representation and linear time-invariant systems.
@@
 It combines the timing dependence modeling capabilities of RNNs [10] with the parallel computing advantages of the Transformer to achieve linear time complexity of long-distance dependence modeling.
-However, the irregular topological characteristics of point cloud data make it difficult for the SSM discrete state transfer mechanism to effectively capture geometric correlations.
-Furthermore, SSM’s unidirectional scanning paradigm is inadequate in modeling the intrinsically bidirectional spatial dependencies embedded within point clouds, particularly in local feature refinement processes where this limitation inevitably induces structural information loss of geometric details.
-To address these limitations, this paper proposes a Geometry-Aware and Cross-Path Feature Hybrid Enhanced Mamba Model (Hybrid-Emba3D), which realizes lightweight local spatial information feature enhancement and feature cross-path fusion, improving the ability to extract local and global information.
+Pointmamba [11] pioneers the application of Mamba models to point cloud processing but neglects local context modeling.
+The irregular topological characteristics of point cloud data make it difficult for the global sequence modeling method of SSM to effectively capture geometric correlations, thereby overlooking local neighborhood relationships and structural details.
+In contrast, Mamba3D [12] catches local features of point clouds through Local Norm Pooling and bidirectional SSM (BiSSM).
+However, recent work E-Mamba [13] adopts local geometry pooling primarily for token reordering, which fails to establish explicit geometric-feature dependencies, limited ability to perceive local structures.
+More critically, the existing Mamba-based point cloud processing methods primarily focus on refining the input scanning strategy of SSMs, that is, reordering or diversifying scanning inputs.
+Although some methods incorporate geometric information, they serve merely to facilitate reordering without directly modeling the relationships between geometric structures and features[14, 15, 16], this may make the model more inclined to learn patterns based on fixed scan order rather than true spatial structure.
+In addition, SSM’s global sequential modeling struggles to directly extract hierarchical features, inevitably leading to deficient local refinement during feature extraction processes.
+To address these limitations, this paper proposes a Geometry-Aware and Cross-Path Feature Hybrid Enhanced Mamba Model (HyMamba).
+Unlike E-Mamba’s pooling strategy, our Geometric-Feature Coupled Pooling (GFCP) injects geometric priors into feature representation without extra parameters, fundamentally differentiating it from existing local pooling methods.
+Specifically, HyMamba realizes: (1) Lightweight local spatial enhancement via geometric-aware feature coupling; (2) Cross-path feature fusion through our Collaborative feature enhancer.
+based on this unique design, our framework achieves state-of-the-art(SOTA) accuracy of 95.99% on ModelNet40, significantly outperforming existing methods and establishing new performance benchmarks as shown in Figure 1.
+Figure 1: Comprehensive comparison of HyMamba with other SOTA models. (All results are based on Self-supervised Pre-training.)
 The main contributions of this work can be summarized as follows:
   • 1.
-    A lightweight feature coupling mechanism based on geometric priors is proposed, which couples geometric feature information between center point features and neighboring point spatial geometric information, enriching local geometric representations.
+    We proposed an efficient point cloud classification model. This model has a strong feature learning ability, achieves extremely high classification performance.
   • 2.
-    Designing a multi-dimensional hybrid feature enhancement module with channel-wise and multi-scale learning capabilities, which enhances context-aware modeling capacity for SSM.
+    We designed the Geometric-Feature Coupled Pooling that dynamically fuses neighbor point geometry with center features via Gaussian spatial weights, enhancing the representational ability of local geometry.
   • 3.
-    The proposed model demonstrates exceptional performance, particularly outperforming existing SOTA models on the ModelNet40 benchmark and establishing new performance benchmarks.
-As shown in Figure 1.
-Figure 1: Comprehensive comparison of Hybrid-Emba3D with other SOTA models
+    We proposed the Collaborative Feature Enhancer that addresses the absence of hierarchical features in global SSM modeling through dual-path hybrid enhancement of channel and spatial features.
+  • 4.
+    The experiment showed that a new record accuracy of 95.99% was achieved on ModelNet40 with 3.95G loating Point Operations (FLOPs), demonstrating excellent efficiency-accuracy balance.
 2 Related work
 2.1 Deep learning for point cloud data
-Effective point cloud classification requires the collaborative modeling of local geometric details and global spatial context.
-Maturana[29] et al.
-proposed VoxNet, which organizes unstructured point clouds into 3D voxel grids and applies conventional 3D convolutional neural networks to capture comprehensive spatial contextual features.
-Based on PointNet, PointNet++[31] employs Farthest Point Sampling (FPS) to construct hierarchical architectures while enhancing localized detail extraction through its local feature aggregation mechanism.
-Its density-adaptive modules effectively mitigate multi-scale learning interference caused by non-uniform point density distributions.
-Subsequent approaches encompass point-based direct convolution[36, 26, 9, 42], Volume-Based convolution [34, 35, 17, 43], and graph-based convolution[41, 28, 40, 24].
-They are all attempts to enhance local feature extraction through spatial transformations or deformable convolutional kernels.
-However, these methodologies fundamentally suffer from the mismatch between fixed receptive fields and irregular point distributions.
-In 2021, PointTransformer[8] was the first to introduce the Transformer architecture, marking a new stage in global context modeling for point cloud processing.
+The efficacy of point cloud classification hinges upon synergistic modeling of local and global contexts.
+Early approaches such as VoxNet [17] and OctNet [18] leverage auxiliary voxel data structures and 3D convolutions to address spatial disorder in point clouds.
+Nevertheless, this reliance on intermediate representations inevitably induces substantial feature degradation.
+To address this limitation, Qi et al.’s PointNet pioneered direct point set processing via shared multi-layer perceptrons (MLP) and symmetric functions.
+Building upon this foundation, PointNet++ [19] enhances local detail extraction by introducing farthest point sampling (FPS) to establish the hierarchical structure.
+Graph convolutional approaches further enhance local geometric feature extraction by establishing connections between neighboring points through directed graphs.
+For example, DGCNN [20] fuses edge features into the graph convolutional operations, consequently enhancing classification accuracy.
+Methodologies including Volume-based 3D convolutions [21, 22, 23], point-based direct convolutions [24, 25, 26, 27], and graph convolutional networks [20, 28, 29, 30] fundamentally rely on local feature aggregation.
+These methodologies fundamentally suffer from the mismatch between fixed receptive fields and irregular point distributions, which makes it difficult to model the long-distance dependency in point cloud data and limits its performance improvement in complex scenes.
+Owing to its excellent global modeling capabilities, Vision Transformer [31] (ViT) has become one of the mainstream architectures in point cloud analysis.
 Based on relative position coding, its self-attention mechanism can adaptively establish geometric correlations between point cloud elements and significantly improve the scene semantic understanding ability through global dependency modeling [32, 33, 34, 35, 36].
-Notably, the architecture’s intrinsic permutation equivariance naturally conforms with the unordered nature of point clouds, eliminating the need for data rearrangement or interpolation required by voxel/grid-based methods, thereby avoiding the information loss caused by three-dimensional coordinate discretization.
 The conflict exists between the sparse distribution characteristics of point cloud data and the dense connectivity assumption underlying standard attention mechanisms, leading to redundant attention weight allocations.
@@
 Such as DSVT [37] with hierarchical feature aggregation demonstrates optimized computation-accuracy balance through adaptive sparse attention patterns; Point-BERT [33] self-supervised pretraining via masked point modeling, effectively reducing annotation dependency; PatchFormer [38] significantly reduces computational complexity through geometrically consistent patch partitioning and streamlined patch-level attention mechanisms.
-While convolution-based architectures exhibit linear computational complexity, their localized receptive fields fundamentally constrain long-range contextual modeling capabilities.
-Conversely, Transformer models achieve global interaction through self-attention mechanisms, yet are inherently limited by quadratic computational complexity.
-In this context, state space models (SSMs) like Mamba demonstrate potential for point cloud processing, achieving linear modeling characteristics in long-sequence modeling that significantly reduce both computational and spatial overhead while preserving global dependency learning.
-2.2 State Space Model
-SSM, a classic method in control theory, achieves dynamic system modeling by constructing the state space.
+In this context, Mamba demonstrates potential for point cloud processing, achieving linear complexity in long-sequence modeling while possessing global learning ability.
+2.2 Selective State Space Model
+SSM, a classic method in control theory, achieves system modeling by constructing the state space.
 As control theory evolved, the limitations of traditional input-output models have gradually emerged.
 SSM establishes a system state space based on input signal characteristic variables, uses state variables to characterize the dynamic evolution of the system fully, and abstracts the system into a mathematical model containing state and output equations.
-The SSM structure are shown in Figure 2.
+Discretize the continuous-time linear time-invariant (LTI) state space model using the zero-order hold (ZOH) method.
+The discrete state and output equations of the discretized SSM are presented in Equation (1), with the corresponding architecture illustrated in Figure 2.
+[…]
+Where: $\bar{A}$, $\bar{B}$, $\bar{C}$, and $\bar{D}$ are the discrete state transition matrix, control input matrix, observation matrix, and feedforward matrix.
+Specifically, $\bar{A}={\exp{(\Delta A)}}$, $\bar{B}={{(\Delta A)}^{-1}}(\bar{A}-I)\Delta B$, $\ Delta$is the discrete sampling interval of ZOH.
+${h}_{k}$ is the current state variable, ${h}_{k-1}$ represents the hidden state of the system at the discrete time point $k-1$, containing historical information up to that time, and $k$ is the discrete time step index.
 Figure 2: Mamba block and SSM structure
-Discretize the continuous-time linear time-invariant (LTI) state space model using the zero-order hold (ZOH) method. The discretized equation of state is given in equation 1.
-${}\left\{\begin{matrix}{{h}_{k}}&=\bar{A}\cdot{{h}_{k-1}}+\bar{B}\cdot{{x}_{k}}\\ {{y}_{k}}&=\bar{C}\cdot{{h}_{k}}+\bar{D}\cdot{{x}_{k}}\\ \end{matrix}\right.$ (1)
-Where: $\bar{A}$ is a discrete state transition matrix, $\bar{B}$ is a discrete control input matrix, $\bar{C}$ is a discrete observation matrix, and $\bar{D}$ is a discrete feedforward matrix.
-${h}_{k}$ is the current state variable, ${h}_{k-1}$ represents the hidden state of the system at the discrete time point $k$-1, containing historical information up to that time, and $k$ is the discrete time step index.
 Traditional SSMs [39, 40, 41]employ static predefined parameters for matrices $\bar{A}$ , $\bar{B}$ , and $\bar{C}$ , resulting in rigid information propagation pathways.
-In contrast, Mamba’s SSM[11] layer implements a selective mechanism that dynamically generates timestep-specific parameters $\Delta$ conditioned on input data.
-This enables adaptive modulation of the discretization processes for B and C matrices, thereby establishing context-sensitive information propagation paths.
-Generate parameters $\Delta$, $\bar{B}$, $\bar{C}$ in real time from the input sequence through linear transformation:
-${{B}_{k}}=\text{Linea}{{\text{r}}_{B}}({{x}_{k}})\in{{\mathbb{R}}^{N\times 1}}$ (2)
-${{C}_{k}}=\text{Linea}{{\text{r}}_{C}}({{x}_{k}})\in{{\mathbb{R}}^{N\times 1}}$ (3)
-${{\Delta}_{k}}=\text{softplus}(\text{Linea}{{\text{r}}_{\Delta}}({{x}_{k}}))\in{{\mathbb{R}}^{+}}$ (4)
-Where: Linear represents linear projection, and $\Delta$ is the step length at which ZOH discretizes the continuous equation.
-Dynamically adjust the discretization process according to the input:
-${{{\bar{A}}}_{k}}=\exp({{\Delta}_{k}}{{A}_{k}})$ (5)
-${{{\bar{B}}}_{k}}={{({{\Delta}_{k}}{{A}_{k}})}^{-1}}(\exp({{\Delta}_{k}}{{A}_{k}})-I){{\Delta}_{k}}{{B}_{k}}$ (6)
-${{{\bar{C}}}_{k}}={{C}_{k}}$ (7)
-Plugging parameter $\bar{A}$, $\bar{B}$, $\bar{C}$ into Equation 1, the selective SSM still follows the LTI property in the continuous time domain, thus allowing efficient GPU-accelerated convolution computing of the discretized state space equation:
-$\bar{K}=(\overline{{{C}_{k}}{{B}_{k}}},{{\overline{{{C}_{k}}{{A}_{k}}B}}_{k}},...,\overline{{{C}_{k}}A}_{k}^{L-1}{{\overline{B}}_{k}})$ (8)
-${{y}_{k}}={{x}_{k}}*\bar{K}$ (9)
-Where: $\bar{K}$ is the global convolution kernel, $L$ is the length of the convolution kernel.
-In summary, the selective SSM’s continuous-time equation is formally consistent with traditional SSM, but content-aware modeling is realized through the dynamic parameterization of the discretization process.
-This design paradigm transcends LTI limitations while inheriting the theory of continuous systems, enabling Mamba to simultaneously attain computational efficiency comparable to RNNs (O(n) complexity) and expressive power equivalent to Transformer architectures.
-PointMamba[21] pioneers the introduction of Mamba architecture into point cloud processing, which significantly improves efficiency while maintaining high accuracy, demonstrating the potential of Mamba in 3D vision tasks.
-3 Hybrid-Emba3D
+In contrast, Mamba’s SSM [9] layer implements a dynamic mechanism that dynamically adaptive modulation of ${B}_{k}$, ${C}_{k}$ and ${\Delta}_{k}$ depending input data:
+${{B}_{k}}={Linear}_{B}({{x}_{k}})$ (2)
+${{C}_{k}}={Linear}_{C}({{x}_{k}})$ (3)
+${{\Delta}_{k}}={softplus}({Linear_{\Delta}}({{x}_{k}}))$ (4)
+Where: $Linear$ represents linear projection.
+Then discretize the dynamic parameters, at this time $\bar{A}_{k}={\exp({{\Delta}_{k}A})}$, ${{{\bar{B}}}_{k}}={{({{\Delta}_{k}}{{A}})}^{-1}}(\bar{A}_{k}-I){{\Delta}_{k}}{{B}_{k}}$, ${{{\bar{C}}}_{k}}={{C}_{k}}$.
+Plugging parameter $\bar{A}_{k}$, $\bar{B}_{k}$, $\bar{C}_{k}$ into equation (1), Selective SSMs effectively transform SSMs into time-varying systems, thereby establishing context-sensitive information propagation paths.
+However, this adaptation renders parallelization through convolution not viable.
+Consequently, Mamba introduces a parallel scan algorithm to achieve efficient parallel computing:
+${{y}_{k}}$ $={\bar{C}_{k}}{{h}_{k}}={\bar{C}_{k}}Scan({{{\bar{A}}}_{k}},{{{\bar{B}}}_{k}}{{x}_{k}})$ (5)
+$={\bar{C}_{k}}\left(\left(\prod\nolimits_{i=1}^{k}{\left[\begin{matrix}{{{\bar{A}}}_{i}}&{{{\bar{B}}}_{i}}{{x}_{i}}\\ 0&1\\ \end{matrix}\right]}\right)\left[\begin{matrix}0\\ 1\\ \end{matrix}\right]\right)$
+Overall, the selective SSM’s continuous-time equation is formally consistent with traditional SSM, but content-aware modeling is realized through the parameter dynamization in the discretization process.
+This design paradigm transcends LTI limitations, enabling Mamba to simultaneously attain computational efficiency comparable to RNNs (O(n) complexity) and global expressive power equivalent to Transformer architectures in long sequence input.
+3 HyMamba
 While Mamba achieves input-dependent parameter adaptation through selective gating mechanisms, its core architecture remains anchored in RNN-style sequential recurrence.
@@
 This work establishes a hybrid framework that effectively enhances the representation performance of the Mamba.
-It integrates the Local geometric pooling (LGP) with geometric feature coupling and a bidirectional SSM enhanced by the collaborative feature enhancer (CoFE-BiSSM).
+It integrates the Geometry-Feature Coupled Pooling (GFCP) and a bidirectional SSM enhanced by the collaborative feature enhancer (CoFE-BiSSM).
 The architecture of the model is shown in Figure 3.
-Figure 3: The architecture of Hybrid-Emba3D.
+Figure 3: The architecture of HyMamba.
 The input point cloud is sampled into local patches via FPS and KNN, and the coordinates and features are aggregated into embedding vectors through lightweight PointNet.
-After incorporating positional encoding and classification tokens, input the Hybrid-Emba3D encoder.
-The encoder core is composed of alternately stacked CoFE-BiSSM and LGP, supplemented by normalization and residual connections.
+After incorporating positional encoding and classification tokens, input the HyMamba encoder.
+The encoder core is composed of alternately stacked CoFE-BiSSM and GFCP, supplemented by normalization and residual connections.
 Task-specific heads finally decode the transformed features for the classification of objects.
-Given an input point cloud $P\in{{\mathbb{R}}^{N\times 3}}$ with $N$ points, the framework first initializes $L$ centroid points via Farthest Point Sampling (FPS) to form a set ${{P}_{\text{center}}}\in{{\mathbb{R}}^{L\times 3}}$.
-For each centroid ${{P}_{\text{center}}}$, the method retrieves its K-Nearest Neighbors (KNN) from the original cloud to construct local patches $x_{p}^{i}\in{{\mathbb{R}}^{K\times 3}}$ serving as structural spatial primitives.
-A lightweight PointNet variant subsequently performs hierarchical feature encoding on each patch, with the resultant local descriptors simultaneously serving as input token sequences for Hybrid-Emba3D architectures.
+Given an input point cloud $P\in{{\mathbb{R}}^{N\times 3}}$ with $N$ points, the framework first initializes $G$ centroid points via Farthest Point Sampling (FPS) to form a set ${{P}_{{center}}}\in{{\mathbb{R}}^{G\times 3}}$.
+For each centroid ${{P}_{{center}}}$, the method retrieves its K-Nearest Neighbors (KNN) from the original cloud to construct local patches $x_{p}^{i}\in{{\mathbb{R}}^{K\times 3}}$ serving as structural spatial primitives.
+A lightweight PointNet variant subsequently performs hierarchical feature encoding on each patch, with the resultant local descriptors simultaneously serving as input token sequences for HyMamba architectures.
 Adhering to the ViT’s [31] work, the architecture embeds a learnable [class] token preceding $L$ local tokens for global representation aggregation.
-Each patch ${{x}_{p}}\in{{\mathbb{R}}^{L\times K\times 3}}$ is projected and mapped into a C-dimensional feature vector ${{W}_{patch}}({{x}_{p}})\in{{R}^{L\times C}}$ as its initial patch embeddings.
-The [class] token is vertically stacked with all point cloud block features, forming a sequence matrix with a shape of ($L$+1)×$C$ ($L$ patch features $+$ one [class]).
+Each patch ${{x}_{p}}\in{{\mathbb{R}}^{G\times K\times 3}}$ is projected and mapped into a C-dimensional feature vector ${{W}_{patch}}({{x}_{p}})\in{{R}^{G\times C}}$ as its initial patch embeddings.
+The [class] token is vertically stacked with all point cloud block features, forming a sequence matrix with a shape of ($G$+1)×$C$ ($G$ patch features $+$ one [class]).
 After superimposing position encoding ${P}_{pos}$, the geometric structure information of the original point cloud is preserved through patch position encoding, providing an initial representation for subsequent hierarchical encoding.
-${{Z}^{0}}=[{{x}_{\text{cls}}}\text{+}P_{\text{pos}}^{(0)},{{W}_{patch}}(x_{p}^{1})+P_{\text{pos}}^{(1)},\cdots,{{W}_{patch}}(x_{p}^{L})+P_{\text{pos}}^{(L)}]$ (10)
+${{Z}^{0}}=[{{x}_{{cls}}}{,}P_{{pos}}^{(0)};{{W}_{patch}}(x_{p}^{1}),P_{{pos}}^{(1)};\cdots;{{W}_{patch}}(x_{p}^{G}),P_{{pos}}^{(G)}]$ (6)
 Where: ${Z}^{0}$ is the initial input layer feature, ${{x}_{cls}}\in{{\mathbb{R}}^{1\times C}}$ is the [class] token, ${{W}_{patch}}(x_{p}^{1}),{{W}_{patch}}(x_{p}^{2})\cdots,{{W}_{patch}}(x_{p}^{L})$ is the concatenated point cloud patch feature processed by the lightweight PointNet variant (${W}_{patch}$), and ${P}_{pos}$ is the positional encoding.
-The encoder employs dual-phase processing (The encoder block is shown in Figure 3 (b)): For the $i$-th layer, input features ${Z}^{i-1}$ undergo normalization before the LGP component executes geometric-aware local feature extraction to generate ${{\hat{Z}}^{i}}$.
+The encoder employs dual-phase processing (The encoder block is shown in Figure 3 (b)): For the $i$-th layer, input features ${Z}^{i-1}$ undergo normalization before the GFCP component executes geometric-aware local feature extraction to generate ${{\hat{Z}}^{i}}$.
 Post residual summation, secondary normalization precedes the CoFE-BiSSM component, which supplants conventional attention mechanisms, enabling dynamic feature enhancement and long-range dependency capture, ultimately outputting the current layer representation ${Z}^{i}$.
@@
 The process is as follows:
-${{{\hat{Z}}}^{i}}=\text{LGP}(\text{Norm}({{Z}^{i-1}}+{{P}_{\text{pos}}}))+{{Z}^{i-1}},$ (11)
-${{{Z}}}^{i}=\text{CoFE-BiSSM}(\text{Norm}({{{\hat{Z}}}^{i}}))+{{{\hat{Z}}}^{i}}$ (12)
+${{{\hat{Z}}}^{i}}={GFCP}({Norm}({{Z}^{i-1}}+{{P}_{{pos}}}))+{{Z}^{i-1}},$ (7)
+${{{Z}}}^{i}={CoFE-BiSSM}({Norm}({{{\hat{Z}}}^{i}}))+{{{\hat{Z}}}^{i}}$ (8)
 The architecture employs task-specific heads (e.g., classification MLPs) for diverse downstream tasks, while integrating learnable positional embeddings within each encoder layer to enhance spatial awareness.
 This structure ingeniously transplants the achievements of the Transformer into the Mamba series.
-3.2 Local geometric pooling with geometry-feature coupling mechanism
+3.2 Geometry-Feature Coupled Pooling
 Local features are crucial for point cloud feature learning.
@@
 Therefore, we have designed a local geometry-feature coupling mechanism, which mainly contains local feature normalization, local geometry patch normalization, and feature coupling operations.
-Figure 4: Illustration of LGP.
-This process selects neighboring points through FPS and KNN, extracts normalized features from the center points, and element-wise multiplication with the broadcast extended local geometric weights to achieve the fusion of features and geometric information, followed by a shared MLP for advanced feature extraction.
-According to Section 3.1 and Figure 4, $L$ local domains with $k$ neighbors are constructed using KNN around each center point.
-The center point of these local domains is ${{P}_{\text{center}}}\in{{\mathbb{R}}^{L\times 3}}$ , and the coordinates of the neighborhood point are ${{k}_{i}}$.
-For the local feature normalization, given the center point feature and its K nearest neighbor feature ${{\tilde{F}}_{K}}\in{{\mathbb{R}}^{L\times K\times C}}$, the relative feature offset $\Delta F={{F}_{K}}-{{F}_{C}}$ is first calculated and channel normalized:
-${{\tilde{F}}_{K}}=\frac{\Delta F}{\sqrt{Var(\Delta F)+\varepsilon}}$ (13)
-The normalized neighbor features and the center point features are concatenated along the channel, and the propagation features can be generated through learning an affine transformation:
+Figure 4: Illustration of GFCP.
+Local neighborhoods are constructed based on KNN, and normalized features and coordinates of neighborhood points are extracted.
+These neighborhood features are concatenated with the central feature and processed through the affine transformation to generate the augmented feature.
+Concurrently, geometric weights are computed based on neighborhood coordinates and broadcast on the channel dimension.
+The features are aggregated after weighted fusion.
+Finally, advanced features are extracted through the shared MLP.
+According to Section 3.1 and Figure 4, $G$ local domains are constructed around each center point.
+The central features and coordinates of these local domains are ${{F}_{c}}\in{{\mathbb{R}}^{G\times C}}$ and ${{P}_{c}}\in{{\mathbb{R}}^{G\times 3}}$.
+The $k$ nearest neighbor points are calculated via Euclidean distance within local regions to establish geometric neighborhoods capturing point cloud spatial topology, with neighborhood features ${{F}_{k}}\in{{\mathbb{R}}^{G\times k\times C}}$ and coordinates ${{P}_{k}}\in{{\mathbb{R}}^{G\times k\times 3}}$.
+For the local feature extension, the relative feature offset $\Delta F={{F}_{K}}-{{F}_{C}}$ is first calculated between neighboring features and central features, and channel normalized:
+${{\tilde{F}}_{K}}=\frac{\Delta F}{\sqrt{Std(\Delta F)+\varepsilon}}$ (9)
+Where: ${{\tilde{F}}_{k}}\in{{\mathbb{R}}^{G\times k\times C}}$ is the normalized neighborhood features.
+The normalized neighborhood features and the center point features are concatenated along the channel, and the Extended features can be generated through learning an affine transformation:
 ${{\hat{F}}_{K}}\in{{\mathbb{R}}^{L\times K\times 2C}}=Concat({{\tilde{F}}_{K}},{{F}_{C}})\odot\gamma+\beta$ (10)
-Where: $\gamma$ and $\beta$ are trainable proportional coefficients and shift vectors, respectively.
-That implements feature-aware adaptive updates by explicitly encoding the center-neighborhood feature relationship.
-For the local geometric patch normalization, firstly compute the mean coordinate $\mu$ of $M$ points within the patch, then convert the coordinates of each point into offsets $\Delta p={{p}_{j}}-u$ relative to the patch centroid to eliminate global translational effects.
-Subsequently, calculate the coordinate standard deviation $\sigma={\sqrt{\sum\nolimits_{x=1}^{M}{||\Delta{{p}_{j}}||_{2}^{2}+\varepsilon}}}/{M}\;$, and normalize the coordinates by $p_{j}^{{}^{\prime}}={\Delta{{p}_{j}}}/{\sigma}\;$ to eliminate local scale differences.
-This process ensures that the geometric encoding achieves translation-scale invariance.
+Where: $\gamma$ and $\beta$ are the trainable scaling coefficient and bias vector for the channel dimension, respectively.
+That implements dynamically adjusts the feature distribution and channel importance by explicitly encoding the center-neighborhood feature relationship and adaptive parameters $\gamma$, $\beta$, thereby achieving feature-aware adaptive updates.
+For local geometric weighting, the generation of geometric weights needs to be based on normalized neighborhood coordinates.
+Firstly, convert the coordinates of the neighborhood point into offsets relative $\Delta P={{P}_{k}}-{{P}_{c}}$ to the central coordinate to eliminate global translational effects.
+Subsequently, calculate the neighborhood coordinate standard deviation, and normalize the neighborhood coordinates to eliminate local scale differences.
 The process can be formalized as:
-$p_{j}^{{}^{\prime}}=\frac{{{p}_{j}}-\mu}{\sqrt{\frac{1}{M}\sum\nolimits_{k=1}^{M}{||{{p}_{k}}-\mu||_{\text{2}}^{\text{2}}+\varepsilon}}}$ (15)
-Where: ${{p}_{j}},{{p}_{k}}\in{{\mathbb{R}}^{3}}$ are 3D coordinates of the $j$-th or $k$-th point in the local geometric patch; $\mu$ is the Centroid of the geometric patch, $\mu={\sum\nolimits_{k=1}^{N}{{{p}_{k}}}}/{M}\;$; $j$, $k$ are point indices, where $j$ is used for calculating the offset of individual points, and $k$ is used for summation traversal over all points; $\varepsilon$ is the numerical stability term.
+${{\tilde{P}}_{k}}=\frac{\Delta P}{\sqrt{Std(\Delta P)+\varepsilon}}$ (11)
+Where: ${{\tilde{P}}_{k}}\in{{\mathbb{R}}^{G\times k\times 3}}$ is the normalized neighborhood coordinates.
 This standardization operation enables local geometric features to have translation invariance and scale invariance, significantly improving the model’s generalization ability to geometric deformations.
 Furthermore, Gaussian spatial weights are constructed based on the Euclidean distances between each point in the local patch and the center point to model the geometric structure correlation:
-$\Delta{{p}_{i}}={{p}_{i}}-{{p}_{c}},$ (16)
-${{w}_{i}}=\exp(-||\Delta{{p}_{i}}|{{|}_{2}}),$ (17)
-$W=Broadcast({{w}_{i}})\in{{\mathbb{R}}^{2C\times L\times K}}$ (18)
-${{F}_{weighted}}={{{\hat{F}}}_{k}}\odot W$ (19)
-Where: ${{p}_{i}},{{p}_{c}}\in{{\mathbb{R}}^{3}}$ represent the coordinates of patch points and the center point; ${{w}_{i}}\in{{\mathbb{R}}^{L\times K\times 1}}$ is a weight coefficient based on distance attenuation, ${{\hat{F}}_{k}}$ is a normalized feature tensor, with a shape of $2C$×$L$×$K$; $\odot$ represents element-wise multiplication; $Broadcast(\cdot)$ extends the weight vector to align with the feature dimension.
-The final feature ${{F}_{agg}}\in{{\mathbb{R}}^{2C\times L}}$ is obtained using the Softmax weighted average to achieve adaptive aggregation of neighborhood features:
-${{F}_{agg}}=\frac{\sum\nolimits_{k=1}^{K}{{{F}_{weighted}}\centerdot\exp({{F}_{weighted}})}}{\sum\nolimits_{k=1}^{K}{\exp({{F}_{weighted}})}}$ (20)
-Finally, channel alignment is completed through the shared MLP.
+${{w}_{i}}=\exp(-||{{{\tilde{P}}}_{k}}|{{|}_{2}}),$ (12)
+${{W}_{geo}}=Broadcast({{w}_{i}})\in{{\mathbb{R}}^{G\times k\times 2C}}$ (13)
+Where: ${{w}_{i}}\in{{\mathbb{R}}^{L\times K\times 1}}$ is a weight coefficient based on distance attenuation to reduce the impact of distant points on extended features; $Broadcast(\cdot)$ expands the weight vector to align with the feature dimension, where all channels within the same neighborhood share identical weights, achieving differentiated weighting exclusively in the spatial dimension.
+Finally, performs element-wise multiplication between geometric weights and expanded features, thereby coupling local features with geometric structural information.
+${{F}_{weighted}}={{\hat{F}}_{k}}\odot{{W}_{geo}}$ (14)
+Where: ${{F}_{weighted}}\in{{\mathbb{R}}^{G\times k\times 2C}}$ is the geometrically weighted feature; $\odot$ represents element-wise multiplication;
+The final local feature ${{F}_{agg}}\in{{\mathbb{R}}^{G\times 2C}}$ is obtained using the nonlinear Softmax-like Pooling to achieve adaptive aggregation of neighborhood features:
+${{F}_{agg}}=\frac{\sum\nolimits_{k=1}^{K}{({{F}_{weight}}\odot\exp({{F}_{weight}}))}}{\sum\nolimits_{k=1}^{K}{\exp({{F}_{weight}})}}$ (15)
+[…]
 This component constructs parameter-free geometric weights using Gaussian kernel functions, leveraging distance-dependent decay characteristics between points to reinforce local structural awareness.
@@
 Through a coupling mechanism between spatial distribution and feature representation, it effectively enhances sensitivity to local structures.
-3.3 Bidirectional SSM with the Collaborative feature enhancer
+3.3 BiSSM with the Collaborative feature enhancer
+Figure 5: Details of Collaborative feature enhancer
 While Mamba achieves theoretical benefits in long-sequence modeling through SSMs, its RNN-based recurrent architecture inherently suffers from constrained effective receptive fields due to error accumulation in sequential processing, exhibiting a deficiency in long-range dependency modeling efficiency compared to Transformer architectures.
@@
 The structure is shown in Figure 5.
-Figure 5: Details of Collaborative feature enhancer
-Given the input tensor $X\in{{\mathbb{R}}^{B\times C\times L}}$ divided into $g$ groups along the channel dimension:
-$X^{\prime}=\operatorname{Reshape}(X,[B\cdot g,\frac{C}{g},L])$ (21)
+Given the input feature $X\in{{\mathbb{R}}^{B\times C\times L}}$ divided into $g$ groups along the channel dimension:
+$X^{\prime}={Reshape}(X,[B\cdot g,\frac{C}{g},L])$ (16)
 Construct parallel dual-path features—the 3×3 convolutional path and the gated normalized path to extract local details and global context:
@@
 Perform cross-channel global feature compression on two paths, compress into a single channel, preserve information in the spatial dimension $L$, and serve as the basis for spatial attention weights:
-$\phi({{X}_{i}})={{F}_{soft}}({{P}_{avg}}({{X}_{i}}))\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (24)
-Where: ${F}_{soft}$ is the row-wise Softmax.
+$\phi({{X}_{i}})={{F}_{soft}}({{P}_{avg}}({{X}_{i}}))\in{{\mathbb{R}}^{(B\cdot g)\times 1\times{}^{C}\!\!\diagup\!\!{}_{g}\;}}$ (19)
+Where: ${F}_{soft}$ is the Softmax along a specific dimension.
 Subsequently, cross-path bidirectional interaction is performed to multiply the compressed feature matrices $\phi({X}_{1})$ and $\phi({X}_{2})$:
@@
 Finally, merge the weights and adjust the output shape:
-$\text{output}^{\prime}=X^{\prime}\odot W\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (28)
-$\text{output}=\operatorname{Reshape}(\text{output}^{\prime},(B,C,L))$ (29)
+${output}^{\prime}=X^{\prime}\odot W\in{{\mathbb{R}}^{(B\cdot g)\times{}^{C}\!\!\diagup\!\!{}_{g}\;\times L}}$ (23)
+${output}={Reshape}({output}^{\prime},(B,C,L))$ (24)
 CoFE employs dual pathways to capture local-global contexts, where a dynamic weight matrix explicitly fuses cross-scale features via bidirectional interaction (${X}_{1}\rightleftarrows{X}_{2}$).
-This interaction computes spatial affinities to eliminate unidirectional bias and adaptively models non-uniform dependencies in long sequences, delivering a lightweight attention enhancement for BiSSM (see Figure 6).
+This interaction computes spatial affinities to eliminate unidirectional bias and adaptively models non-uniform dependencies in long sequences, delivering a lightweight attention enhancement for SSM (see Figure 6).
 Figure 6: The structure of CoFE-BiSSM
 The unidirectional nature of Mamba restricts global dependency capture in point cloud processing.
-Unlike Vision Mamba’s[53] token-level horizontal flipping, a reverse path through vertical flipping of feature channels is constructed.
-The resulting BiSSM module is formulated as:
-$F^{\prime}=\text{CoFE}(Con{{v}_{1\times 1}}(F))$ (30)
-[…]
+Unlike Vision Mamba’s [42] token-level horizontal flipping, Mamba3D [12] constructs a reverse path through vertical flipping of feature channels.
+The resulting CoFE-BiSSM module is formulated as:
+$F^{\prime}={Conv}_{1\times 1}({CoFE}(F))$ (25)
+${CoFE-BiSSM}=Linear(F^{\prime}+Fli{{p}_{C}}(F^{\prime}))$ (26)
 Where:$Fli{{p}_{C}}(\cdot)$ is the Channel Flip operation.
 This approach eliminates reliance on point cloud token order, enabling the model to focus on inherent distribution patterns of feature vectors rather than artificially constructed sequential orders.
-While Mamba’s core mechanism captures long-range dependencies through parameterized state transition equations, it demonstrates insufficient sensitivity to local transient features.
-The CoFE component complements the SSM framework by directly modeling inter-position relationships.
+While Mamba’s core mechanism captures long-range dependencies through state transition equations’ parameter dynamization, it demonstrates insufficient sensitivity to local transient features.
+The CoFE component complements the SSMs framework by directly modeling inter-position relationships.
 The ${X}_{2}$ pathway reduces the limitations of SSMs in capturing local details through localized context aggregation.
-Subsequent ${X}_{1}\rightleftarrows{X}_{2}$ interaction facilitates complementary integration of feature correlation and provides more discriminative input to SSMs, enabling SSMs to dynamically perceive global contextual patterns while effectively modeling complex sequential dependencies.
+Subsequent ${X}_{1}\rightleftarrows{X}_{2}$ interaction facilitates complementary integration of feature correlation and provides more discriminative input to SSMs, enhancing BiSSM to dynamically perceive global contextual patterns while effectively modeling complex sequential dependencies.
 4 Experiments
@@
 Unless explicitly specified, consistent training parameters are maintained between scratch training and fine-tuning.
-The Hybrid-Emba3D classification task header utilizes a 3-layer MLP classifier (256→256→$C$ channels).
+The HyMamba classification task header utilizes a 3-layer MLP classifier (256→256→$C$ channels).
 In the ModelNet40 experiment, adopting the standard division (9843 train/2468 test), sampling 1024 points, and using scale&translation augmentation, while ScanObjectNN followed the original train/test split, sampling 2048 points, and using the rotation augmentation.
@@
 Table 1: Experimental results of ModelNet40 and ScanObjectiNN.
-Compared with methods of different architectures: 3D understanding architectures, Transformer-based architectures, •Diffusion-driven frameworks, Mamba-based architectures.
-* denotes the application of voting strategy.
-The results of other models are sourced from publicly published papers.
 Methods  Reference Pre-training strategy ModelNet40 1k P ScanObjectNN P (M) F (G)
-OBJ_BG OBJ_ONL PB_T50_RS
+OBJ_BG OBJ_ONLY PB_T50_RS
 Supervised Learning Only
@@
 Mamba3D [12] ACM MM24 × 93.4 92.94 92.08 91.81 16.9 3.9
-Hybrid-Emba3D × 94.0 93.63 92.25 91.05 16.96 3.95
+HyMamba × 94.0 93.63 92.25 91.05 16.96 3.95
 With Self-supervised Pre-training
 Transformer [6] NeurIPS 17 OcCo 92.1 84.85 85.54 78.79 22.1 4.8
-MaskPoint[23] CVPR 22 MaskPoint 93.8 89.30 88.10 84.30 22.1 4.8
+MaskPoint [51] CVPR 22 - 93.8 89.30 88.10 84.30 22.1 4.8
 Point-BERT* [33] CVPR 22 IDPT 93.4 88.12 88.30 83.69 22.1+1.7 4.8
 Point-MAE* [32] ECCV 22 IDPT 94.4 91.22 90.02 84.94 22.1+1.7 4.8
-PointGPT-S*[5] NeurIPS 23 94.0 91.6 90.0 86.9 29.2 5.7
+PointGPT-S* [52] NeurIPS 23 - 94.0 91.6 90.0 86.9 29.2 5.7
 Point-FMAE* [53] AAAI 24 Point-M2AE 95.15 95.18 93.29 90.22 27.4 3.6
-PointDif[52] CVPR 24 PointDif - 93.29 91.91 87.61 - -
+PointDif [54] CVPR 24 - - 93.29 91.91 87.61 - -
 Pointmamba [11] NeurIPS 24 Point-MAE 93.6 94.32 92.60 89.31 12.3 3.6
 Mamba3D* [12] ACM MM24 Point-MAE 95.1 95.18 92.60 93.34 16.9 3.9
-Hybrid-Emba3D Point-MAE 95.58+1.58 93.80+0.27 93.12+0.87 93.09+2.04 16.96 3.95
-Hybrid-Emba3D* Point-MAE 95.99+1.99 95.18+1.38 93.80+1.55 94.34+3.29 16.96 3.95
-Comparison with existing SOTA methods: Hybrid-Emba3D achieves SOTA performance across multiple benchmarks.
+E-Mamba [13] NeuCom 25 Point-MAE 94.7 94.32 92.94 91.98 13.78 7.49
+HyMamba Point-MAE 95.58+1.58 93.80+0.27 93.12+0.87 93.09+2.04 16.96 3.95
+HyMamba* Point-MAE 95.99+1.99 95.18+1.38 93.80+1.55 94.34+3.29 16.96 3.95
+  • 1
+    Compared with methods of different architectures: 3D understanding architectures, Transformer-based architectures, •Diffusion-driven frameworks, Mamba-based architectures.
+* denotes the application of voting strategy.
+Comparison with existing SOTA methods: HyMamba achieves SOTA performance across multiple benchmarks.
 Without pretraining, the model attains 94.0% accuracy on ModelNet40, surpassing the comparable Mamba3D by 0.6%, while achieving the highest OBJ_BG accuracy of 93.63%.
-When employing Point-MAE pretraining, Hybrid-Emba3D breaks the 95.58% accuracy barrier on ModelNet40, further reaching 95.99% with the voting strategy, establishing a new record for this benchmark.
-Hybrid-Emba3D outperforms SOTA Transformer counterparts (PointGPT-S, Point-FMAE) by 1.99% and 0.84% in accuracy, while reducing Parameters by 41.9% and 39.1%, respectively.
+When employing Point-MAE pretraining, HyMamba breaks the 95.58% accuracy barrier on ModelNet40, further reaching 95.99% with the voting strategy, establishing a new record for this benchmark.
+HyMamba outperforms SOTA Transformer counterparts (PointGPT-S, Point-FMAE) by 1.99% and 0.84% in accuracy, while reducing Parameters by 41.9% and 39.1%, respectively.
 Compared to similar Mamba-based models (Pointmamba, Mamba3D), it achieves 0.89% and 1.0% accuracy gains on ModelNet40 and ScanObjectNN-PB_T50_RS (the most challenging benchmark) with only exchanged 0.06M Parameters and 0.05G FLOPs.
+Notably, HyMamba demonstrates clear advantages over the latest E-Mamba in both accuracy (+0.88% on ModelNet40, +1.11% on PB_T50_RS) and computational efficiency (3.95G vs 7.49G FLOPs) when using pretrained weights.
+Although E-Mamba introduces geometric information in the process of reordering, it is fundamentally different from the explicit and direct geometric feature coupling of HyMamba.
+With pre-trained weights from ShapeNetCore providing a rich initial representation, this distinction becomes critical.
+GFCP can leverage this rich foundation to more accurately simulate complex local geometry, ultimately unlocking the full potential of the model to achieve significant performance improvements.
 Feature visualization: High-dimensional features are reduced via t-SNE to a 2D projection distribution to visualize the clustering effect of features.
 As shown in Figure 7.
-On ModelNet40, the feature distribution is relatively scattered, and there are ambiguous inter-class boundaries, indicating that the Hybrid Emba3D needs to improve its ability to distinguish specific categories.
+On ModelNet40, the feature distribution is relatively scattered, and there are ambiguous inter-class boundaries, indicating that the HyMamba needs to improve its ability to distinguish specific categories.
 While OBJ_BG and PB_T50_RS demonstrate prominent clustering tendencies with minor inter-class overlaps, OBJ_ONLY achieves superior class separation.
 The model maintains robust clustering performance under PB_T50_RS interference scenarios, confirming its resilience to environmental disturbances.
-Figure 7: Details of Collaborative feature enhancer
-Figure 8: Details of Collaborative feature enhancer
+Figure 7: t-SNE diagram of HyMamba on the ModelNet40 and ScanobjectNN.
 Analysis of analysis category confusion: Overall performance is good on ModelNet40 (see Table 1 and Figure 8), but similar geometric structures and overlapping local features result in a specific category-15 (dresser) accuracy of only 35% (consistent with t-SNE), and its confusion rates with category-26 (piano) and category-37 (tv_stand) are 35% and 25%, respectively.
+Figure 8: Confusion matrix of HyMamba on the ModelNet40 and ScanobjectNN.
 Comparing OBJ_ONLY and OBJ_BG, the presence of background increases the misjudgment rate of locally structurally similar objects.
-For example, the accuracy of category-0 (bag) decreases from 100% to 88%, and the confusion rate of category-5 (chair) and 9 (shelf) reaches 13%.
-Conversely, it may also reduce the misclassifications rate caused by feature overlap, such as the accuracy rate of category-10 and 12 increased from 86% and 75% to 95% and 88%, respectively, and the misjudgment rate of category-9 (shelf) and 10 (table), category-11 (bed) and category-12 (pillow) drops from 9% and 12% to 0%.
+For example, the accuracy of category-0 (bag) decreases from 100% to 88%, and the confusion rate of category-5 (chair) and 9 (shelves) reaches 13%.
+Conversely, it may also reduce the misclassifications rate caused by feature overlap, such as the accuracy rate of category-10 and 12 increased from 86% and 75% to 95% and 88%, respectively, and the misjudgment rate of category-9 (shelves) and 10 (table), category-11 (bed) and category-12 (pillow) drops from 9% and 12% to 0%.
 Under interference (PB_T50_RS), the misclassification distribution has shifted, meaning misjudged samples are redistributed among error categories while maintaining a stable overall misjudgment rate.
+4.3 Ablation study
+To validate the complementary nature and distinct performance boundaries of core components GFCP and CoFEBiSSM, across diverse scenarios, we conduct ablation studies on the HyMamba benchmarked on 3D point cloud classification tasks.
+The results as shown in Table 2
 Table 2: Ablation studies on ModelNet40 and ScanObjectNN.
-LGP CoFE-BiSSM ModelNet40 1k P ScanObjectNN P (M) F (G)
+GFCP CoFE-BiSSM ModelNet40 1k P ScanObjectNN P (M) F (G)
 OBJ_BG OBJ_ONL PB_T50_RS
@@
 94.00 93.63 92.25 91.05 16.96 3.95
+GFCP: GFCP is a core component focusing on local feature extraction.
+This component captures subtle geometric variations and topological structures on object surfaces by constructing non-linear transformation mechanisms within local neighborhoods.
+When solely activated, GFCP achieves 94.00% accuracy on ModelNet40, demonstrating its strong geometric feature representation in ideal, background-free scenarios.
+Furthermore, it achieves accuracy gains of +0.97% (OBJ_BG), +1.04% (OBJ_ONLY), and +0.35% (PB_T50_RS) compared to the deactivated baseline, further validating its capability to resist background clutter and minor deformations through local geometric modeling.
+It is commendable that it introduces zero additional parameters (0M) and computational costs (0G FLOPs), highlighting superior lightweight design advantages.
+CoFE-BiSSM: CoFE-BiSSM enhances long-range dependencies through multi-scale feature fusion and bidirectional modeling.
+It achieves 92.05% accuracy on OBJ_BG, slightly below GFCP’s 92.57%, but with excellent efficiency (0.03M Parameters/0.09G FLOPs).
+In the ModelNet40 (93.56%) and OBJ-ONLY (90.02%) scenarios without background interference, the performance is relatively limited, while the rotation disturbance sensitivity on PB_T50_RS (90.32%) reveals insufficient local geometric modeling.
+Experiments demonstrate its effectiveness in global context reasoning for complex background tasks, but it requires integration with local feature extraction modules to address geometric sensitivity.
+Joint effect: Component collaboration yields significant performance gains.
+At the peak of 94.00% (+0.61%) in ModelNet40, all subsets of ScanObjectNN surpassed the single component configuration comprehensively, with OBJ-BG improving by 2.03%, OBJ-ONLY improving by 2.41%, and PB_T50_S improving by 0.73%.
+They achieve feature complementarity with minimal overhead.
+In summary, GFCP and CoFE-BiSSM focus on local geometric and global multi-scale modeling, respectively.
+While their strengths are scenario-specific, joint deployment enables complementary features and enhances generalization performance on different datasets.
+4.4 Few-shot Learning
+Datasets and Settings: ModelNetFewShot is the ModelNet40-based 3D few-shot benchmark for evaluating classification algorithms’ generalization with limited samples.
+It adopts N-way K-shot tasks, randomly selecting N classes (K training samples per class) and testing on the remaining instances.
+Results report mean accuracy ± standard deviation across 10 independent trials.
+The results are shown in Table 3.
 Table 3: The experimental results of ModelNetFewShot. The results of other models are sourced from publicly published papers.
@@
 DGCNN+CrossPoint [55] CVPR 2022 92.5±3.0 94.9±2.1 83.6±5.3 87.9±4.2
-Hybrid-Emba3D 90.5±3.8 96.0±3.2 86.3±5.1 92.0±3.4
+HyMamba 90.5±3.8 96.0±3.2 86.3±5.1 92.0±3.4
 With Self-supervised Pre-training
@@
 Pointmamba [11] NeurIPS 24 96.9±2.3 99.0±1.1 93.0±4.0 95.6±2.8
-Hybrid-Emba3D 93.3±5.0 98.9±1.3 90.4±5.7 95.6±3.9
-4.3 Ablation study
-To validate the complementary nature and distinct performance boundaries of core components LGP and CoFEBiSSM, across diverse scenarios, we conduct ablation studies on the Hybrid-Emba3D benchmarked on 3D point cloud classification tasks.
-The results as shown in Table 2
-LGP: LGP is a core component focusing on local feature extraction.
-This component captures subtle geometric variations and topological structures on object surfaces by constructing non-linear transformation mechanisms within local neighborhoods.
-When solely activated, LGP achieves 94.00% accuracy on ModelNet40, demonstrating its strong geometric feature representation in ideal, background-free scenarios.
-Furthermore, it achieves accuracy gains of +0.97% (OBJ_BG), +1.04% (OBJ_ONLY), and +0.35% (PB_T50_RS) compared to the deactivated baseline, further validating its capability to resist background clutter and minor deformations through local geometric modeling.
-It is commendable that it introduces zero additional parameters (0M) and computational costs (0G FLOPs), highlighting superior lightweight design advantages.
-CoFE-BiSSM: CoFE-BiSSM enhances long-range dependencies through multi-scale feature fusion and bidirectional modeling.
-It achieves 92.05% accuracy on OBJ_BG, slightly below LGP’s 92.57%, but with excellent efficiency (0.03M Parameters/0.09G FLOPs).
-In the ModelNet40 (93.56%) and OBJ-ONLY (90.02%) scenarios without background interference, the performance is relatively limited, while the rotation disturbance sensitivity on PB_T50_RS (90.32%) reveals insufficient local geometric modeling.
-Experiments demonstrate its effectiveness in global context reasoning for complex background tasks, but it requires integration with local feature extraction modules to address geometric sensitivity.
-Joint effect: Component collaboration yields significant performance gains.
-At the peak of 94.00% (+0.61%) in ModelNet40, all subsets of ScanObjectNN surpassed the single component configuration comprehensively, with OBJ-BG improving by 2.03%, OBJ-ONLY improving by 2.41%, and PB_T50_S improving by 0.73%.
-They achieve feature complementarity with minimal overhead.
-In summary, LGP and CoFE-BiSSM focus on local geometric and global multi-scale modeling, respectively.
-While their strengths are scenario-specific, joint deployment enables complementary features and enhances generalization performance on different datasets.
-4.4 Few-shot Learning
-Datasets and Settings: ModelNetFewShot is the ModelNet40-based 3D few-shot benchmark for evaluating classification algorithms’ generalization with limited samples.
-It adopts N-way K-shot tasks, randomly selecting N classes (K training samples per class) and testing on the remaining instances.
-Results report mean accuracy ± standard deviation across 10 independent trials.
-The results are shown in Table 3.
-Results: Without pretraining, Hybrid-Emba3D achieves 90.5% on 5-Way 10-Shot, surpassing Transformer (87.8±5.2) but lagging behind DGCNN-CrossPoint (92.5%).
+HyMamba 93.3±5.0 98.9±1.3 90.4±5.7 95.6±3.9
+Results: Without pretraining, HyMamba achieves 90.5% on 5-Way 10-Shot, surpassing Transformer (87.8±5.2) but lagging behind DGCNN-CrossPoint (92.5%).
 Scaling to 20-shot boosts accuracy by 6% (96.0%).
 For 10-Way tasks, it outperforms DGCNN-CrossPoint.
-With self-supervised pretraining, it is worth noting that Hybrid Emba3D outperforms PointGPT-S (98.6%) in 5-Way 20 Shot, achieving a quasi SOTA performance of 98.9%.
+With self-supervised pretraining, it is worth noting that HyMamba outperforms PointGPT-S (98.6%) in 5-Way 20 Shot, achieving a quasi SOTA performance of 98.9%.
 However, the advantage has not been generalized to other settings.
@@
 5 Conclusion
-In this paper, we present Hybrid-Emba3D, a novel mamba-based Architecture for point cloud classification task, designing two key components: the geometry-feature coupling mechanism that aggregates geometric information from neighboring points around each centroid and central feature to enhance local geometric representation, and A multidimensional feature hybrid enhancer employing dual-path Architecture to strengthen global context modeling for BiSSM.
+In this paper, we present HyMamba, a novel mamba-based Architecture for the point cloud classification task, designing two key components: the geometry-feature coupling mechanism that aggregates geometric information from neighboring points around each centroid and central feature to enhance local geometric representation, and a multidimensional feature hybrid enhancer employing the dual-path architecture to strengthen global context modeling for BiSSM.
 Notably, the total computational cost with the two components is very small.
-Especially in the operations of coupling local geometric information into central features in a parameter-free manner, incurring no additional parameters, that is completely driven by local neighborhoods’ intrinsic geometric relationships.
-Extensive experiments demonstrate Hybrid-Emba3D’s superior performance across classification benchmarks, particularly achieving new SOTA accuracy on ModelNet40 at the cost of minimal parameters.
+Especially in the operations of coupling local geometric information into central features in a parameter-free manner, which is completely driven by local neighborhoods’ intrinsic geometric relationships.
+Extensive experiments demonstrate HyMamba’s superior performance across classification benchmarks, particularly achieving new SOTA accuracy on ModelNet40 at the cost of minimal parameters.
 These also show that the Mamba architecture has potential, yet insufficient exploration in point cloud processing.
+Future research could dynamically adjust the point cloud serialization strategy or SSM scanning strategy based on the relationships between points, so that the scanning path can reflect the spatial relationships between points to compensate for its lack of directness in modeling 3D spatial structures.
+Additionally, integrating CNN architectures or optimizing SSM’s long-sequence processing could enhance the representation of hierarchical local features.
+Acknowledgements
+All authors thank the 512 Lab and 513 Lab of the School of Weapon Science and Technology at Xi’an Technological University.
+CRediT authorship contribution statement
+Bin Liu: Conceptualization, Data curation, Formal analysis, Methodology, Project administration, Resources, Software, Writing – original draft.
+Chunyang Wang: Funding acquisition, Supervision.
+Xuelian Liu: Visualization, Validation.
+Bo Xiao: Investigation, Supervision.
+Guan Xi: Supervision,.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Data availability
+Data will be made available on request.